# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 5.2.1 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)>0$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ \frac{-1}{Q(z)}&0\end{bmatrix}$ $z\in\Gamma_{a_{-}}^{b}$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z){\rm e}^{n\Omega_{2}}\\ \frac{-{\rm e}^{-n\Omega_{2}}}{Q(z)}&0\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$M_{{}_{Q}}(z_{+})$ $=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ \frac{-{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1},$ (5.57)
with $\varkappa$ as in (5.39) and $Q$ as in (5.3). Furthermore the following local behaviours hold
$M_{{}_{Q}}(z)$ $=\mathbf{1}+\mathcal{O}(z^{-1})$ $|z|\to\infty$ (5.58)
$M_{{}_{Q}}(z)$ $=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\ \ \ q=1,b,a_{\pm}$ $z\to q.$ (5.59)
To further normalize the problem we need to construct a Szegö function, along the line of what we already have done in the case of the outside of $EoT$.
The Szegö function.
Let $R(z)$ denote the radical function
$R(z)=\sqrt{(z-1)(z-b)(z-a_{+})(z-a_{-})}=\sqrt{z^{4}+(A-2)z^{3}+\left(\frac{s^{2}}{4}-A+1\right)z^{2}-\frac{s^{2}}{2}z+\frac{s^{2}}{4}},$ (5.60)
@@
$S(z)=R(z)\left(\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}+\int_{a_{-}}^{b}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}+\int_{b}^{a_{+}}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}\right).$ (5.61)
The constant $\nu$ is chosen so that the function $S(z)$ is bounded as $|z|\to\infty$. We can actually simplify (5.61) quite a bit; consider, for example the integral
$I_{1}:=\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2i\pi}.$ (5.62)
Since, along $\Gamma_{b}^{1}$ we have $Q(w_{+})=Q(w_{-}){\rm e}^{2i\pi{{\rho}}}$ and $R(w_{+})=-R(w_{-})$ we can convert $I_{1}$ into
$I_{1}=\frac{1}{2}\int_{b}^{1}\frac{\ln Q(w_{+}){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}+\frac{1}{2}\int_{1}^{b}\frac{\ln Q(w_{-}){\mathrm{d}}w}{R(w_{-})(w-z)2i\pi}+\int_{b}^{1}\frac{i\pi{{\rho}}+\ln\varkappa-\nu{\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}.$ (5.63)
Along the same lines, denoting the other two integrals in (5.61) by $I_{2},I_{3}$ we have
$I_{2}=\frac{1}{2}\int_{a_{-}}^{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}+\frac{1}{2}\int^{a_{-}}_{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{-})(w-z)2i\pi}$ (5.64)
$I_{3}=\frac{1}{2}\int^{a_{+}}_{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}+\frac{1}{2}\int_{a_{+}}^{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{-})(w-z)2i\pi}.$ (5.65)
Adding $I_{1}+I_{2}+I_{3}$ we observe that the integrals involving $\ln Q(w)$ form a closed loop from $b$ that goes around the whole $\Gamma_{m}$ in the clockwise direction without intersecting the cut of $Q(z)$. Using (5.3) we observe that we have
$\ln Q(z_{+})=\ln Q(z_{-})+\begin{cases}2i\pi{K}&z\in(-\infty,0)\\ 2i\pi{\rho}&z\in(0,1).\end{cases}$ (5.66)
By using Cauchy’s theorem we conclude that
$S(z)$ $=\frac{1}{2}\ln Q(z)+R(z)\left(\int_{-\infty}^{0}\frac{-{K}{\mathrm{d}}w}{2R(w)(w-z)}-\frac{1}{2}\int_{b}^{0}\frac{{{\rho}}{\mathrm{d}}w}{R(w)(w-z)}+\int_{b}^{1}\frac{(i\pi{{\rho}}+\ln\varkappa-\nu){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}\right).$ (5.67)
Observe that the branch cut of $\ln Q(z)$ on $(-\infty,0)$ is compensated by the branch cut of the first integral in the above expression, so that $S(z)$ is actually analytic across the ray $(-\infty,0)$. From (5.67) it is then simple to ascertain that to make $S(z)$ bounded at infinity we need to choose the value of $\nu$ as follows (see (5.39) for $\varkappa$)
$\nu=i\pi{{\rho}}\left(1-\frac{\int_{b}^{0}\frac{{\mathrm{d}}w}{R(w)}}{\int_{b}^{1}\frac{{\mathrm{d}}w}{R(w_{+})}}\right)+\ln\varkappa.$ (5.68)
The expression (5.61) makes it clear, through the use of the Sokhostki–Plemelji formula, that the following properties of the Szegö function hold:
Proposition 5.2 (Szegö function for $\mathrm{Re}\,(s)>0$)
@@
$S(z_{+})+S(z_{-})=\left\{\begin{array}[]{lc}\ln\left(\varkappa Q(z_{+})\right)-\nu&z\in\Gamma_{b}^{1}\\ \ln Q(z)&z\in\Gamma_{a_{-}}^{b}\cup\Gamma_{b}^{a_{+}}.\end{array}\right.$ (5.71)
Furthermore $S(z)$ is bounded near $z=b,a_{\pm}$ and near $z=1$ it has the behaviour ([30] Ch. 5)
$S(z)=-\frac{{\rho}}{2}\ln(z-1)+\mathcal{O}(1).$ (5.72)
Alternative description/derivation of the Szegö function.
Proceeding from the desired properties in Proposition (5.2) and differentiating we obtain
$S^{\prime}(z_{+})+S^{\prime}(z_{-})=\frac{Q^{\prime}(z)}{Q(z)},\ \ \ \ z\in\Gamma_{m}.$ (5.73)
@@
where $\eta(z)$ must be a function analytic off $\Gamma_{m}$ with $\eta(z_{+})+\eta(z_{-})=0$ on $\Gamma_{m}$ and such as to cancel out the singularities of $\frac{Q^{\prime}}{Q}$ in (5.74) outside of $\Gamma_{m}$. Since $\frac{Q^{\prime}}{Q}=\frac{{\rho}}{z-1}+\frac{{K}-{\rho}}{z}$ we see that
$\eta(z)=\frac{{\rho}-{K}}{2}\frac{R(0)}{zR(z)}-\frac{C}{R(z)},$ (5.75)
which cancels the pole at $z=0$ in the expression (5.74). Here $R$ is given by (5.60). The constant $C$ is determined by the condition that $S(z_{+})+S(z_{-})=\ln Q(z)$ on the arcs $\Gamma_{b}^{a_{\pm}}$, which implies that integral of $\eta$ along the contour $\widetilde{\gamma}$ on the Riemann surface of $W^{2}=R(z)^{2}$ (see Fig. 9, left pane) vanishes. Specifically it is given by:
@@
Using the explicit expression of $R$ in (5.60) we note that $R(0)=-\frac{s}{2}$ and can simplify (5.76) to
$C=C_{+}=\frac{({K}-{\rho})}{4}\frac{s\oint_{\widetilde{\gamma}_{+}}\frac{{\mathrm{d}}z}{z\,W}}{\int_{b}^{1}\frac{{\mathrm{d}}z}{R(z_{+})}},$ (5.77)
where we have used the fact that $\oint_{\widetilde{\gamma}}\frac{{\mathrm{d}}z}{R(z)}=-\int_{b}^{1}\frac{{\mathrm{d}}z}{R(z_{+})}$ (see Fig. 9, left pane).
When $\mathrm{Re}\,(s)<0$ instead we have to ask that the antiderivative of (5.74) is continuous on ${\mathbb{C}}\setminus\Gamma_{m}$ and then the contour $\widetilde{\gamma}$ is different. In this case then the constant $C$ is given by
@@
We observe that, in fact, the contours $\widetilde{\gamma}_{\pm}$ are homologically equivalent on the Riemann surface punctured at the two points $z=0$ on both sheets. However the definition of the domain of $R(z)$ has been chosen with different branch cuts. Keeping this in mind, we can use for $C$ the expression (5.78) in both cases with the understanding that in the case $\mathrm{Re}\,(s)>0$ the indication of the boundary value $R(z_{+})$ is irrelevant.
@@
$S^{\prime}(0)=-\frac{{\rho}}{2}\frac{R^{\prime}(0)}{R(0)}-\frac{C_{\pm}}{R(0)}=\frac{{\rho}}{2}+\frac{{({K}-{\rho})}}{2}\frac{\oint_{a_{+}}^{a_{-}}\frac{{\mathrm{d}}z}{z\,R(z_{+})}}{\oint_{a_{+}}^{a_{-}}\frac{{\mathrm{d}}z}{R(z_{+})}}.$ (5.79)
Figure 9: The alternative computation of the Szego function and the contour $\widetilde{\gamma}$. For $\mathrm{Re}\,(s)>0$ in the left pane, for $\mathrm{Re}\,(s)<0$ in the right pane. The dashed part of the contour $\widetilde{\gamma}$ corresponds to the second sheet of the radical function $R(z)$ (the one where $R(z)\simeq-z^{2}$ at infinity). The two contours are homologically equivalent.
With the Szegö function at hand we can normalize the RHP 5.8. We define a new model problem
@@
The effect of this normalization is to turn the jump matrices (5.57) into constant ones as detailed in the following problem
Problem 5.9
Find a matrix–valued function $M(z)$ analytic in ${\mathbb{C}}\setminus\Gamma_{m}$ such that
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ $z\in\Gamma_{a_{-}}^{b}$
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{2}}\\ {-{\rm e}^{-n\Omega_{2}}}&0\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$M(z_{+})$ $=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{1}+\nu}\\ -{\rm e}^{-n\Omega_{1}-\nu}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1}$ (5.81)
@@
$M(z)$ $=\mathbf{1}+\mathcal{O}(z^{-1})$ $|z|\to\infty$ (5.82)
$M(z)$ $=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\ \ \ q=1,b,a_{\pm}$ $z\to q.$ (5.83)
The solution of the RHP 5.9 is relatively standard and it requires the use of the Jacobi theta function.
Solution of the RHP 5.9.
With the same definition of $R$ in (5.60) we also define
@@
$h(z_{+})$ $=ih(z_{-})$ $z\in\Gamma_{a_{-}}^{b}$ (5.85)
$h(z_{+})$ $=-ih(z_{-})$ $z\in\Gamma_{b}^{1}\cup\Gamma_{b}^{a_{+}}.$ (5.86)
We also need the Abel map: let
@@
$\mathfrak{u}(z):=\int_{a_{-}}^{z}\frac{{\mathrm{d}}w}{2\omega_{1}R(w)},\ \ \ \ \tau:=\frac{\omega_{2}}{\omega_{1}},$ (5.88)
where the contour of integration runs in the simply connected domain ${\mathbb{C}}\setminus\Gamma$. A direct inspection reveals
Lemma 5.1 (Properties of the Abel map)
@@
$\mathfrak{u}(z_{+})=-\mathfrak{u}(z_{-})+\left\{\begin{array}[]{lc}0&z\in\Gamma_{a_{-}}^{b}\\ -1&z\in\Gamma_{b}^{a_{+}}\\ -\tau&z\in\Gamma_{b}^{1}.\end{array}\right.$ (5.92)
Let $\vartheta(u;\tau)$ be the Riemann (Jacobi) theta function (also denoted $\theta_{4}$ in DLMF 20.2.4 but with a different normalization for $u$) :
$\vartheta(u;\tau):=\sum_{n\in{\mathbb{Z}}}{\rm e}^{in^{2}\pi\tau+2i\pi nu}.$ (5.93)
@@
$\vartheta(u+k+\ell\tau;\tau)={\rm e}^{-2i\pi\ell u-i\pi\ell^{2}\tau}\vartheta(u;\tau).$ (5.94)
Consider the following two row-vectors;
$\boldsymbol{\phi}(z;\infty)$ $=[\phi_{1}(z;\infty),\phi_{2}(z;\infty)],$
$\phi_{1}(z;\infty)$ $=\frac{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{i\pi K\mathfrak{u}(z)}$
$\phi_{2}(z;\infty)$ $=\frac{-i\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-i\pi K\mathfrak{u}(z)}$ (5.95)
@@
$\psi_{1}(z;\infty)$ $=\frac{-i\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{i\pi K\mathfrak{u}(z)}$
$\psi_{2}(z;\infty)$ $=\frac{-\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-i\pi K\mathfrak{u}(z)}.$ (5.96)
Furthermore all entries are bounded by (5.83) by the very definition of $h(z)$ (5.84) and the fact that $\vartheta$ is an entire function with zeros only at the half-period.
Let us investigate the behaviour of $\boldsymbol{\phi},\boldsymbol{\psi}$ near $z=\infty$. Observe that $\phi_{2}$ and $\psi_{1}$ tend to zero as $z\to\infty$ (due to the factor $h(z)=\mathcal{O}(z^{-1})$), while $\phi_{1},\psi_{2}$ have to be computed using l’Hopital’s rule given that the $\vartheta$ in the denominator also tends to zero.

### v3

@@ 5.2.1 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)>0}$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ \displaystyle\frac{-1}{Q(z)}&0\end{bmatrix},\qquad z\in\Gamma_{a_{-}}^{b},$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z){\rm e}^{n\Omega_{2}}\\ \displaystyle\frac{-{\rm e}^{-n\Omega_{2}}}{Q(z)}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&\varkappa Q(z_{+}){\rm e}^{n\Omega_{1}}\\ \displaystyle\frac{-{\rm e}^{-n\Omega_{1}}}{\varkappa Q(z_{+})}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1},$ (5.18)
with $\varkappa$ as in (5.10) and $Q$ as in (5.1). Furthermore, the following local behaviours hold:
$M_{{}_{Q}}(z)=\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr),\qquad|z|\to\infty,$
$M_{{}_{Q}}(z)=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\qquad q=1,b,a_{\pm},\qquad z\to q.$
To further normalize the problem, we need to construct a Szegő function, along the line of what we already have done in the case of the outside of $\rm EoT$}.
The Szegő function. Let $R(z)$ denote the radical function
$R(z)$ $=\sqrt{(z-1)(z-b)(z-a_{+})(z-a_{-})}$
$=\sqrt{z^{4}+(A-2)z^{3}+\left(\frac{s^{2}}{4}-A+1\right)z^{2}-\frac{s^{2}}{2}z+\frac{s^{2}}{4}},$ (5.19)
@@
$S(z)={}$ $R(z)\left(\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}+\int_{a_{-}}^{b}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}\right$.
$\left.+\int_{b}^{a_{+}}\frac{\ln Q(w)}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}\right).$ (5.20)
The constant $\nu$ is chosen so that the function $S(z)$ is bounded as $|z|\to\infty$. We can actually simplify (5.20) quite a bit. Consider, for example, the integral
$I_{1}:=\int_{b}^{1}\frac{\ln Q(w_{+})+\ln\varkappa-\nu}{R(w_{+})(w-z)}\frac{{\mathrm{d}}w}{2{\rm i}\pi}.$
Since, along $\Gamma_{b}^{1}$ we have $Q(w_{+})=Q(w_{-}){\rm e}^{2{\rm i}\pi{{\rho}}}$ and $R(w_{+})=-R(w_{-})$, we can convert $I_{1}$ into
$I_{1}=\frac{1}{2}\int_{b}^{1}\frac{\ln Q(w_{+}){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}+\frac{1}{2}\int_{1}^{b}\frac{\ln Q(w_{-}){\mathrm{d}}w}{R(w_{-})(w-z)2{\rm i}\pi}+\int_{b}^{1}\frac{{\rm i}\pi{{\rho}}+\ln\varkappa-\nu{\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}.$
Along the same lines, denoting the other two integrals in (5.20) by $I_{2}$, $I_{3}$, we have
$I_{2}=\frac{1}{2}\int_{a_{-}}^{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}+\frac{1}{2}\int^{a_{-}}_{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{-})(w-z)2{\rm i}\pi},$
$I_{3}=\frac{1}{2}\int^{a_{+}}_{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}+\frac{1}{2}\int_{a_{+}}^{b}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{-})(w-z)2{\rm i}\pi}.$
Adding $I_{1}+I_{2}+I_{3}$, we observe that the integrals involving $\ln Q(w)$ form a closed loop from $b$ that goes around the whole $\Gamma_{m}$ in the clockwise direction without intersecting the cut of $Q(z)$. Using (5.1), we observe that we have
$\ln Q(z_{+})=\ln Q(z_{-})+\begin{cases}2{\rm i}\pi{K},&z\in(-\infty,0),\\ 2{\rm i}\pi{\rho},&z\in(0,1).\end{cases}$
By using Cauchy’s theorem, we conclude that
$S(z)={}$ $\frac{1}{2}\ln Q(z)+R(z)\left(\int_{-\infty}^{0}\frac{-{K}{\mathrm{d}}w}{2R(w)(w-z)}\right$.
$\left.-\frac{1}{2}\int_{b}^{0}\frac{{{\rho}}{\mathrm{d}}w}{R(w)(w-z)}+\int_{b}^{1}\frac{({\rm i}\pi{{\rho}}+\ln\varkappa-\nu){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}\right).$ (5.21)
Observe that the branch cut of $\ln Q(z)$ on $(-\infty,0)$ is compensated by the branch cut of the first integral in the above expression, so that $S(z)$ is actually analytic across the ray $(-\infty,0)$. From (5.21), it is then simple to ascertain that to make $S(z)$ bounded at infinity we need to choose the value of $\nu$ as follows (see (5.10) for $\varkappa$):
$\nu={\rm i}\pi{{\rho}}\left(1-\frac{\int_{b}^{0}\frac{{\mathrm{d}}w}{R(w)}}{\int_{b}^{1}\frac{{\mathrm{d}}w}{R(w_{+})}}\right)+\ln\varkappa.$ (5.22)
The expression (5.20) makes it clear, through the use of the Sokhostki–Plemelji formula, that the following properties of the Szegő function hold.
Proposition 5.14 (Szegő function for $\operatorname{Re}(s)>0$).
@@
$S(z_{+})+S(z_{-})=\begin{cases}\ln\left(\varkappa Q(z_{+})\right)-\nu,&z\in\Gamma_{b}^{1},\\ \ln Q(z),&z\in\Gamma_{a_{-}}^{b}\cup\Gamma_{b}^{a_{+}}.\end{cases}$
Furthermore, $S(z)$ is bounded near $z=b,a_{\pm}$ and near $z=1$ it has the behaviour [29, Chapter 5]
$S(z)=-\frac{{\rho}}{2}\ln(z-1)+\mathcal{O}(1).$
Alternative description/derivation of the Szegő function. Proceeding from the desired properties in Proposition (5.14) and differentiating, we obtain
$S^{\prime}(z_{+})+S^{\prime}(z_{-})=\frac{Q^{\prime}(z)}{Q(z)},\qquad z\in\Gamma_{m}.$
@@
where $\eta(z)$ must be a function analytic off $\Gamma_{m}$ with $\eta(z_{+})+\eta(z_{-})=0$ on $\Gamma_{m}$ and such as to cancel out the singularities of $\frac{Q^{\prime}}{Q}$ in (5.23) outside of $\Gamma_{m}$. Since $\frac{Q^{\prime}}{Q}=\frac{{\rho}}{z-1}+\frac{{K}-{\rho}}{z}$, we see that
$\eta(z)=\frac{{\rho}-{K}}{2}\frac{R(0)}{zR(z)}-\frac{C}{R(z)},$
which cancels the pole at $z=0$ in the expression (5.23). Here $R$ is given by (5.19). The constant $C$ is determined by the condition that $S(z_{+})+S(z_{-})=\ln Q(z)$ on the arcs $\Gamma_{b}^{a_{\pm}}$, which implies that integral of $\eta$ along the contour $\widetilde{\gamma}$ on the Riemann surface of $W^{2}=R(z)^{2}$ (see Figure 9, left pane) vanishes. Specifically, it is given by
@@
Using the explicit expression of $R$ in (5.19), we note that $R(0)=-\frac{s}{2}$ and can simplify (5.24) to
$C=C_{+}=\frac{({K}-{\rho})}{4}\frac{s\oint_{\widetilde{\gamma}_{+}}\frac{{\mathrm{d}}z}{zW}}{\int_{b}^{1}\frac{{\mathrm{d}}z}{R(z_{+})}},$
where we have used the fact that
$\oint_{\widetilde{\gamma}}\frac{{\mathrm{d}}z}{R(z)}=-\int_{b}^{1}\frac{{\mathrm{d}}z}{R(z_{+})}$
(see Figure 9, left pane).
When $\operatorname{Re}(s)<0$ instead, we have to ask that the antiderivative of (5.23) is continuous on ${\mathbb{C}}\setminus\Gamma_{m}$ and then the contour $\widetilde{\gamma}$ is different. In this case then the constant $C$ is given by
@@
We observe that, in fact, the contours $\widetilde{\gamma}_{\pm}$ are homologically equivalent on the Riemann surface punctured at the two points $z=0$ on both sheets. However, the definition of the domain of $R(z)$ has been chosen with different branch cuts. Keeping this in mind, we can use for $C$ the expression (5.25) in both cases with the understanding that in the case $\operatorname{Re}(s)>0$ the indication of the boundary value $R(z_{+})$ is irrelevant.
@@
$S^{\prime}(0)=-\frac{{\rho}}{2}\frac{R^{\prime}(0)}{R(0)}-\frac{C_{\pm}}{R(0)}=\frac{{\rho}}{2}+\frac{{({K}-{\rho})}}{2}\frac{\oint_{a_{+}}^{a_{-}}\frac{{\mathrm{d}}z}{zR(z_{+})}}{\oint_{a_{+}}^{a_{-}}\frac{{\mathrm{d}}z}{R(z_{+})}}.$
Figure 9: The alternative computation of the Szegő function and the contour $\widetilde{\gamma}$. For $\operatorname{Re}(s)>0$ in the left pane, for $\operatorname{Re}(s)<0$ in the right pane. The dashed part of the contour $\widetilde{\gamma}$ corresponds to the second sheet of the radical function $R(z)$ (the one where $R(z)\simeq-z^{2}$ at infinity). The two contours are homologically equivalent.
With the Szegő function at hand, we can normalize the Riemann–Hilbert Problem 5.13. We define a new model problem
@@
The effect of this normalization is to turn the jump matrices (5.18) into constant ones as detailed in the following problem.
Problem 5.15.
Find a matrix-valued function $M(z)$ analytic in ${\mathbb{C}}\setminus\Gamma_{m}$ such that
$M(z_{+})=M(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\qquad z\in\Gamma_{a_{-}}^{b},$
$M(z_{+})=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{2}}\\ {-{\rm e}^{-n\Omega_{2}}}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$M(z_{+})=M(z_{-})\begin{bmatrix}0&{\rm e}^{n\Omega_{1}+\nu}\\ -{\rm e}^{-n\Omega_{1}-\nu}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1}$ (5.27)
@@
$M(z)=\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr),\qquad|z|\to\infty,$ (5.28)
$M(z)=\mathcal{O}\left(\frac{1}{(z-q)^{\frac{1}{4}}}\right),\qquad q=1,b,a_{\pm},\quad z\to q.$ (5.29)
The solution of the Riemann–Hilbert Problem 5.15 is relatively standard and it requires the use of the Jacobi theta function.
Solution of the Riemann–Hilbert Problem 5.15. With the same definition of $R$ in (5.19), we also define
@@
$h(z_{+})={\rm i}h(z_{-}),\qquad z\in\Gamma_{a_{-}}^{b},\qquad h(z_{+})=-{\rm i}h(z_{-}),\qquad z\in\Gamma_{b}^{1}\cup\Gamma_{b}^{a_{+}}.$ (5.31)
We also need the Abel map. Let
@@
$\mathfrak{u}(z):=\int_{a_{-}}^{z}\frac{{\mathrm{d}}w}{2\omega_{1}R(w)},\qquad\tau:=\frac{\omega_{2}}{\omega_{1}},$ (5.33)
where the contour of integration runs in the simply connected domain ${\mathbb{C}}\setminus\Gamma$. A direct inspection reveals the following.
Lemma 5.16 (properties of the Abel map).
@@
$\mathfrak{u}(z_{+})=-\mathfrak{u}(z_{-})+\begin{cases}\phantom{-}0,&z\in\Gamma_{a_{-}}^{b},\\ -1,&z\in\Gamma_{b}^{a_{+}},\\ -\tau,&z\in\Gamma_{b}^{1}.\end{cases}$
Let $\vartheta(u;\tau)$ be the Riemann (Jacobi) theta function (also denoted $\theta_{4}$ in DLMF 20.2.4 but with a different normalization for $u$)
$\vartheta(u;\tau):=\sum_{n\in{\mathbb{Z}}}{\rm e}^{{\rm i}n^{2}\pi\tau+2{\rm i}\pi nu}.$
@@
$\vartheta(u+k+\ell\tau;\tau)={\rm e}^{-2{\rm i}\pi\ell u-{\rm i}\pi\ell^{2}\tau}\vartheta(u;\tau).$ (5.34)
Consider the following two row-vectors:
$\boldsymbol{\phi}(z;\infty)=[\phi_{1}(z;\infty),\phi_{2}(z;\infty)],\qquad\phi_{1}(z;\infty)=\frac{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{{\rm i}\pi K\mathfrak{u}(z)},$
$\phi_{2}(z;\infty)=\frac{-{\rm i}\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-{\rm i}\pi K\mathfrak{u}(z)}$ (5.35)
@@
$\psi_{1}(z;\infty)=\frac{-{\rm i}\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{{\rm i}\pi K\mathfrak{u}(z)},$
$\psi_{2}(z;\infty)=\frac{-\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-{\rm i}\pi K\mathfrak{u}(z)}.$ (5.36)
Furthermore, all entries are bounded by (5.29) by the very definition of $h(z)$ (5.30) and the fact that $\vartheta$ is an entire function with zeros only at the half-period.
Let us investigate the behaviour of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ near $z=\infty$. Observe that $\phi_{2}$ and $\psi_{1}$ tend to zero as $z\to\infty$} (due to the factor $h(z)=\mathcal{O}\bigl(z^{-1}\bigr)$), while $\phi_{1}$, $\psi_{2}$ have to be computed using l’Hopital’s rule given that the $\vartheta$ in the denominator also tends to zero.